### v1

# Like Father, Like Son: Kinship-Aware Preference Mapping (KARMA) for Automatic Alignment in Large Language Models

Abstract
Recent advancements in Large Language Model (LLM) alignment have sought to mitigate the cost of human annotations by leveraging pretrained models to generate preference data. However, existing methods often compare responses from models with substantially different capabilities, yielding superficial distinctions that fail to provide meaningful guidance on what constitutes a superior response. To address this limitation, we propose Kinship-Aware pReference MApping (KARMA), a novel framework that systematically pairs responses from models with comparable competencies. By constraining preference comparisons to outputs of similar complexity and quality, KARMA enhances the informativeness of preference data and improves the granularity of alignment signals. Empirical evaluations demonstrate that our kinship-aware approach leads to more consistent and interpretable alignment outcomes, ultimately facilitating a more principled and reliable pathway for aligning LLM behavior with human preferences.
Like Father, Like Son: Kinship-Aware Preference Mapping (KARMA)
for Automatic Alignment in Large Language Models
Jeesu Jung1, Chanjun Park2, Sangkeun Jung1††thanks: Corresponding author 1Chungnam National University, 2Korea University jisu.jung5@gmail.com, bcj1210@korea.ac.kr, hugmanskj@gmail.com
1 Introduction
Aligning Large Language Models (LLMs) with human preferences is a fundamental challenge in artificial intelligence, where the effectiveness of alignment hinges on the quality and specificity of the preference data used to guide model outputs Shen et al. . Early alignment methodologies primarily relied on human-annotated datasets with binary preference labels, wherein explicit human judgments directed models toward generating more desirable responses Ouyang et al. (2022a). While these methods have played a critical role in improving LLM behavior, their dependence on extensive human labor imposes significant scalability constraints.
To address this limitation, recent research has explored automated preference data generation techniques, such as Reinforcement Learning from AI Feedback (RLAIF) Lee et al. . These methods leverage pretrained models to automatically generate preference pairs, significantly reducing the need for human annotation. However, the quality of the resulting preference data remains a major concern. If responses within a preference pair exhibit substantial disparities in quality, complexity, or style, the preference signal may become trivial, capturing only superficial contrasts rather than informative distinctions that refine model behavior.
We define this issue as the Response Quality Gap, which refers to the disparity in coherence, complexity, and reasoning quality between candidate responses. A large Response Quality Gap weakens preference signals, leading to misaligned or less informative training objectives. This problem is particularly pronounced in large-scale, automatically generated datasets, where manually curating response pairs to minimize the Response Quality Gap is infeasible due to the massive volume of data.
A promising solution lies in considering the origin of the responses being compared—specifically, the LLM that generated them. Since a model’s reasoning ability, commonsense knowledge, and instruction-following capabilities are inherently shaped by its training data and architecture, the quality of its responses is largely determined by these characteristics. Consequently, comparing responses from models with vastly different capabilities introduces inconsistencies in preference data, making alignment objectives less reliable.
To formalize this idea, we introduce the notion of Kinship among LLMs, which quantifies the similarity in reasoning ability, commonsense understanding, and instruction adherence among models. Measured using standard benchmarks, kinship provides a principled way to assess whether two models generate responses of comparable complexity and quality. Building on this insight, we propose the Kinship-Aware pReference MApping (KARMA) framework, which systematically pairs responses based on model kinship. Unlike conventional approaches that indiscriminately compare outputs from models of varying capabilities, KARMA ensures that preference pairs are drawn from models with similar competencies, thereby narrowing the Response Quality Gap. This results in preference data that captures more subtle and meaningful distinctions, leading to higher-quality alignment signals.
The effectiveness of KARMA is illustrated in Figure 1, which provides an overview of the framework, and Figure 2, which details its response evaluation process. By fostering more structured and granular comparisons, KARMA improves the coherence and interpretability of preference data, shifting alignment objectives from filtering out trivial differences to refining preference signals through comparability and consistency. Our empirical results demonstrate that this kinship-aware approach significantly outperforms prior methods such as RLAIF, offering a more robust and scalable solution for aligning LLM behavior with human preferences.
2 Kinship-Aware pReference MApping (KARMA) Framework
Recent approaches to LLM alignment have reduced the reliance on human annotation through automated preference data generation. However, these methods often neglect the foundational influence of the Origin Model. For example, a model trained using Llama’s methodology is going to produce outputs that fall within the distribution of Llama’s training data. Just as a son inherits traits from his father, a model’s quality directly influences the quality of its outputs. when preference pairs are derived from models with markedly disparate capabilities, the resulting comparisons yield only trivial distinctions.
To address this, the proposed Kinship-Aware pReference MApping (KARMA) framework evaluates kinship based on the performance of origin models on selected benchmark metrics. With kinship as a guiding principle for structuring binarized data, the framework systematically generates preference data for alignment tuning.
2.1 Definition of Kinship
In the context of binarized alignment data mapping, the similarity between the responses serves as a fundamental determinant of mapping efficacy. When data pairs are closely aligned, they provide more nuanced and meaningful guidance signals for alignment tuning. Building upon this principle of response similarity, KARMA introduces the concept of kinship among models.
The kinship is measured by evaluating the capabilities of origin models. We measure the benchmarks as the origin model capability. We define a set of candidate origin models, , each evaluated according to multiple dimensions of competency. Let be a set of benchmark tasks. Each benchmark provides an evaluation metric that quantifies the performance of model . This evaluation may include measures of factual correctness, reasoning, and adherence to instructions.
For each pair of models , we define a kinship score , which reflects the similarity of their capabilities:
where is a similarity function. Higher values of indicate closer competency alignment between and . The exact functional form of can vary depending on the chosen benchmarks and may incorporate multiple factors (denoted collectively as , , and ) to minimize bias and ensure comprehensive evaluation.
Descriptive Multiple-Choice Instruction Following IF evalZhou et al. (2023) Knowledge MMLUHendrycks et al. (2020) MMLU-proWang et al. (2024) Reasoning ARC-EasyClark et al. (2018) ARC-ChallengeClark et al. (2018) HellaswagZellers et al. (2019)
2.2 Stage 1: Benchmark Evaluation for Model Competency
To identify which models are suitable candidates for kinship-based pairing, we employ a series of benchmarks that meet three criteria: (1) they are widely validated in prior research, (2) they cover both general knowledge and domain-specific tasks, and (3) they provide well-defined quantitative metrics. A summary of selected benchmarks is provided in Table 1.
Evaluating models on these benchmarks yields performance scores that serve as the basis for computing . The degree of kinship among models varies with the choice of benchmarks. To avoid relying solely on log-probabilities or overly simplistic metrics, we adopt prompt-based evaluations that require each model to produce explicit responses. Appendix C presents the benchmark performance of all models used for kinship calculation, along with example calculations.
The prompt formats are adapted from templates in MMLU-pro Wang et al. (2024) and FLAN Longpre et al. (2023), ensuring compatibility with each model’s input structure and recommended usage patterns. The prompt template can be found in Appendix B.
2.3 Stage 2: Kinship-Based Response Mapping
Once we identify a subset of models whose pairwise kinship scores exceed a predetermined threshold (e.g., ), we set at 0.1, as cosine similarity values below this threshold indicate a lack of meaningful similarity between models. we focus on collecting responses from these closely matched models. Kinship serves as an indicator of model’s ability and similarity. While various criteria can be applied to measure this, we utilize the models’ capabilities and performance similarities that can be estimated through benchmarks. We have defined kinship across three distinct cases:
: Mapping the two origin models with the highest average benchmark scores.
: Mapping the two origin models with the highest cosine similarity between their benchmark scores.
: Mapping the model with the highest benchmark score to the most similar model based on cosine similarity.
Figure 3 shows the each kinship definition and calculation process.
To generate preference data from these kinship relationships, we convert model responses into binary preference pairs using a consistent metric based on normalized benchmark performance. This approach entirely eliminates the need for instance-wise annotations, as it operates at the dataset level, transforming the entire instruction-response dataset into preference data without requiring instance-wise annotations from either models or humans. By leveraging aggregate performance metrics rather than per-instance comparisons, we maintain consistency while removing computational overhead and human cost.
2.4 Stage 3: Automatic Preference Integration with Alignment Processes
To integrate kinship-based mapped data into existing alignment tuning, the data requires a preference-setting process. Accordingly, we construct our kinship-based preference estimation on the principle that higher-performing models tend to produce more reliable outputs.
To quantify this, we calculate a normalized benchmark score that aggregates multiple benchmark performances into a single value. For a model , this score is computed as:
For a pair of kinship-aligned models , we obtain a binary preference label:
Our data maintains the Response Quality Gap through kinship-based model comparisons and can establish preferences via without additional annotation. This provides high-quality training signals, which can be leveraged through reinforcement learning or supervised fine-tuning.
In contrast to existing approaches (e.g., RLAIF Lee et al. ), KARMA’s focus on model kinship produces preference pairs that yield more nuanced insights. This shift in perspective—from relying on comparisons of models with vastly different capabilities to leveraging finely matched pairs—provides a more parsimonious and theoretically grounded method for constructing alignment-relevant preference data.
Baseline Ours RLAIF Total 0.31 0.31 0.29 0.29 MMLU 0.36 0.33 0.34 0.30 MMLU-Pro 0.15 0.15 0.15 0.14 ARC-Easy 0.40 0.40 0.36 0.41 ARC-Challenge 0.40 0.37 0.33 0.39 IFeval 0.28 0.30 0.29 0.23
3 Experiments
In this paper, we propose the KARMA framework. To validate the functionality of this framework and the characteristics of Alignment Tuning, we conducted experiments to address the following Research Questions (RQs):
RQ1: Does Model Kinship represent the models’ similarity?
RQ2: Does Model Kinship affect the binarization process?
RQ3: Does Model Kinship have a greater influence than model quality, as verified through benchmark evaluations?
RQ4: What benchmark represent the model capability: knowledge, instruction-following, and reasoning?
3.1 Alignment Dataset
For preference binarization, we utilized the Ultrafeedback dataset Cui et al. (2023), which comprises responses generated by a diverse set of language models. The dataset includes outputs from both commercial and open-source models, ensuring a broad representation of model capabilities. The commercial models incorporated in the dataset include GPT-4OpenAI (2023), GPT-3.5 TurboOuyang et al. (2022a), and BardWaisberg et al. (2024). Additionally, several models from the Llama family were included, such as Llama-2 (7B, 13B, and 70B)-chatTouvron et al. (2023), UltraLM-13BCui et al. (2024), WizardLM (7B, 13B, and 70B)Xu et al. (2023), Vicuna-33BZheng et al. (2024), and Alpaca-7BTaori et al. (2023). Beyond the Llama-based architectures, the dataset also features responses from other notable models, including Falcon-40B-instructAlmazrouei et al. (2023), MPT-30B-chatTeam (2023), StarChat-BetaTunstall et al. (2023), and Pythia-12BBiderman et al. (2023).
Although the Ultrafeedback dataset contains results from UltraLM-65BCui et al. (2024), its performance could not be accurately assessed. To maintain the reliability of our evaluation, we excluded these results from the final dataset composition.
3.2 Models
To assess the effectiveness of the newly binarized dataset constructed using the KARMA framework, we fine-tuned and evaluated multiple instruction-following models. The selected models for evaluation include Llama-3.1-(3B, 8B)-InstructDubey et al. (2024), representing Llama-based architectures. Additionally, we included Qwen2.5-(3B, 8B)-InstructYang et al. (2024), enabling a comparative analysis across different model families. This evaluation setup ensures a comprehensive assessment of the impact of kinship-aware preference mapping on alignment performance.
3.3 Training Algorithm for Alignment Tuning
We tested two different learning method for alignment tuning using the KARMA Framework:
Supervised Fine-Tuning (SFT)
Direct Preference Optimization (DPO) Rafailov et al. (2024)
The implementation details are provided in Appendix A
4 Experimental Results
4.1 RQ1: Dataset Reconstruction using KARMA
The proposed KARMA framework enables the application of a unified Total Ranking map across the entire dataset, facilitating a structured and consistent preference mapping process. Utilizing this reconstructed preference dataset, we conducted model training while ensuring that each critic’s binarized response was systematically incorporated. The effectiveness of this reconstructed dataset was assessed by evaluating the trained models on standardized benchmarks, including MMLU and IFeval. The detailed results of these evaluations are presented in Appendix C.
For evaluation, we employed prediction-based assessment methodologies across all benchmark tasks. Specifically, for ARC, MMLU, and MMLU-Pro, we adapted the MMLU-Pro evaluation framework, modifying only the multiple-choice options to align with our dataset. For IFeval, we leveraged its native evaluation framework to ensure consistency in assessment.
To analyze the relationships between models, we computed kinship scores by normalizing evaluation results across six benchmark tasks. Figure 4 presents a Principal Component Analysis (PCA)Maćkiewicz and Ratajczak (1993) visualization of these relationships, illustrating distinct clustering patterns among models. Generally, smaller-scale models tend to cluster in the first quadrant, models with closer kinship relationships in the second quadrant, while Llama-based models and models exceeding 10B parameters are predominantly distributed in the third and fourth quadrants. This distribution underscores the critical role of model training data, training algorithms, and scale in determining kinship relationships among models.
These findings highlight the effectiveness of the KARMA framework in reconstructing preference datasets, providing a more structured and informative approach to model alignment.
4.2 RQ2: Performance of KARMA Binarization
To assess the effectiveness of KARMA binarization, we examined whether model performance can serve as an indicator of data quality. KARMA performs binarization in two sequential steps: first, it defines preference pairs based on kinship between models, and second, it ranks these pairs according to their relative performance. Through this process, we aimed to determine whether response similarity—estimated through model similarity—plays a more significant role in preference modeling than simply prioritizing higher-performing models by consistently positioning them as dominant within each pair.
The results of this comparison are presented in Figure 5. Across all benchmark tasks, the kinship-based scoring metric, , demonstrates performance levels nearly equivalent to the existing instance-wise RLAIF method, with a minimal deviation of only 0.008. Notably, for MMLU-Pro and IFeval, KARMA-based binarization even surpasses the RLAIF baseline by approximately 0.01, indicating that structured preference mapping via kinship can yield competitive or superior alignment outcomes. The detailed results, including average scores for each methodology, are summarized in Appendix D.
4.3 RQ3: Model Size and Kinship
The impact of different kinship approaches varies significantly with model size, influencing how models learn from preference data. To investigate this relationship, we conducted experiments using models from the same family but with different parameter counts, specifically comparing small models (3B parameters) and larger models (7-8B parameters). Our analysis reveals a clear pattern in how model size determines the optimal kinship strategy.
Smaller models (3B) exhibit superior performance when trained using the algorithm, which prioritizes learning from response quality within the binarized dataset. In contrast, larger models (7-8B) achieve better results with the algorithm, suggesting that response similarity becomes increasingly important as model size grows.
This trend indicates that model size fundamentally influences how different models leverage preference data. While smaller models benefit more from explicit learning based on absolute quality differences, larger models demonstrate greater sensitivity to the nuanced relationships between similar responses in the training data. This relationship is visualized in Figure 6, illustrating the distinct learning behaviors observed across different model scales.
4.4 RQ4: Model Capability and Kinship Patterns
Models exhibit distinct similarity patterns based on their capabilities, with these patterns varying across different evaluation metrics. Notably, model kinship is influenced by factors such as model scale and training methodology, leading to variations in clustering behavior across different capability dimensions.
Analysis Framework
To systematically investigate these relationships, we analyzed model similarities across three key capability dimensions:
Knowledge-Based Tasks: Unlike reasoning capability, clustering in instruction-following tasks is more strongly aligned with model families rather than size. This indicates that training methodology and architectural choices exert a greater influence on instruction adherence.
Reasoning Capability: Models tend to cluster primarily based on parameter count, suggesting that model size plays a dominant role in shaping reasoning performance.
Instruction-Following (IF) Tasks: A hierarchical influence pattern emerges, where:
At the initial hierarchy, the model family is the primary determinant.
At the higher hierarchy, the model size becomes a stronger predictor of performance.
Figure 7 visualizes these relationships through dendrograms, illustrating the hierarchical clustering patterns that emerge across different capability dimensions.
5 Conclusion
This paper introduces KARMA, a kinship-based framework for generating high-quality preference data for language model alignment without requiring instance-wise annotation. KARMA achieves performance comparable to RLAIF while offering greater computational efficiency and surpassing baselines on MMLU-Pro and IFeval, demonstrating the effectiveness of kinship-aware preference mapping. Our findings show that model size significantly affects the optimal kinship strategy. Smaller models (3B) benefit from prioritizing response quality, while larger models (7-8B) perform better when leveraging response similarity. Additionally, kinship patterns vary across tasks, with reasoning clustering by model size, instruction-following by model family, and knowledge-based tasks exhibiting a hierarchical structure. These results highlight the need for size-adaptive preference learning strategies and demonstrate that high-quality preference data can be generated without costly manual annotation. Future work will extend KARMA to cross-family model comparisons and further explore the relationship between model kinship and human preference judgments.
6 Related Work and Background
6.1 Alignment Tuning
Alignment tuning has become a central focus in enhancing LLMs to meet user expectations and ethical standardsKumar et al. (2024). Various preference-based learning techniques, particularly reinforcement learning methods such as Proximal Policy Optimization (PPO)Schulman et al. (2017), Direct Preference Optimization (DPO)Rafailov et al. (2024), and Optimized Reinforcement Preference Optimization (ORPO)Hong et al. (2024), have been developed to facilitate this tuning process. These methods rely on preference datasets, which typically contain pairs of responses generated by models based on given instructions, with each pair ranked according to human or model-based evaluations. Prominent alignment datasets like UltrafeedbackCui et al. (2023), which gathers extensive human feedback, and HH-RLHFBai et al. (2022), which uses human-annotated preferences, provide foundational resources for alignment. To minimize the reliance on labor-intensive data curation, recent automated approaches have been introduced to filter or regenerate preference data based on specific criteria, promoting data consistency and scalability. However, such approaches often lack nuanced control over data quality, as they fail to consider the role of the origin model’s capabilities in shaping alignment effectiveness Shi et al. (2024).
6.2 Limitations in Existing Preference Data Approaches
Effective preference data construction requires a clear, rigorous set of criteria to ensure alignment quality across generated response pairs. Common criteria, including Reasoning, Truthfulness, and Instruction-Following, guide the selection of data that aligns with key ethical and functional standardsCui et al. (2023); Bai et al. (2022). High-performing models, capable of producing responses that meet these standards, are often evaluated using benchmarks like ARCClark et al. (2018), MMLUHendrycks et al. (2020), and the Instruction-Following evalZhou et al. (2023) suite, which assess a model’s factual accuracy, reasoning ability, and compliance with instructions. While these benchmarks provide a foundation for selecting responses that promote alignment goals, existing methods typically overlook the impact of variation in model capabilities on alignment consistency.
The KARMA Framework addresses this limitation by introducing a kinship-aware approach that emphasizes model compatibility in preference data selection. By curating preference pairs based on model kinship—aligning models with similar core competencies in knowledge, reasoning, and instruction-following—the KARMA Framework enhances alignment coherence and ensures stable, high-quality data. Further, its curriculum-based selection process sequences preference data from simpler to more complex distinctions, optimizing the alignment process for improved consistency and scalability.
Limitation
In this paper, we demonstrated that KARMA can achieve performance comparable to existing RLAIF without requiring annotations for SFT and DPO. However, due to resource and time constraints, we were unable to validate our approach across a broader range of alignment tuning techniques. Further evaluation is needed for methods.
Additionally, our evaluation primarily focused on fundamental abilities such as Reasoning, Knowledge, and Instruction-Following. However, we did not assess KARMA’s performance on other important values, including factuality and ethical standard. Future work should incorporate evaluations reflecting these aspects.
Appendix A Implementation Details for Model Training
The model was trained for a single epoch using bfloat16 (bf16) quantization, which optimizes memory efficiency while preserving numerical precision. The training configuration incorporated the following hyperparameters: a per-device batch size of 6, gradient accumulation steps set to 4, and a learning rate of , with 500 warm-up steps to facilitate stable convergence.
Training was conducted on an L40 4-GPU setup, leveraging an optimized deep learning framework to enhance computational efficiency. The training pipeline focused on performance optimization through checkpointing and logging, without intermediate model evaluation during training.
Appendix B Prompt template for Evaluation Benchmarks
To determine kinship relationships between models, this paper evaluates benchmark performance and compares the similarity of these numerical results. The benchmark performance of all 17 models considered in this study is presented in Table 8. For the start and end tokens, the template recommended in the paper was used.
Appendix C Evaluation Origin Model Capabilities for Kinship
To determine kinship relationships between models, this paper evaluates benchmark performance and compares the similarity of these numerical results. The benchmark performance of all 17 models considered in this study is presented in Table 3. The actual calculation example for and can be found in Figure 9.
Due to the unavailability of Bard, its performance metrics have been substituted with those of Gemini-1.5-Flash.
Model IFeval MMLU-STEM MMLU-pro Hellaswag ARC-easy ARC-Challenge Average Score GPT-4 85.37 86.40 0.64 95.30 96.63 96.40 0.99 GPT-3.5-turbo 72.54 70.00 0.38 85.00 92.80 83.02 0.77 bard 89.33 71.80 0.59 84.70 84.43 77.13 0.86 Llama-2-7b-chat 25.19 53.10 0.20 67.50 72.14 54.61 0.41 Llama-2-13b-chat 24.82 57.80 0.25 71.20 72.05 58.02 0.45 Llama-2-70b-chat 24.07 68.90 0.38 78.10 82.20 67.66 0.58 UltraLM-13b 54.92 49.58 0.19 54.00 57.58 48.04 0.40 WizardLM-7b 45.83 42.50 0.18 77.70 39.48 34.90 0.34 WizardLM-13b 33.92 52.30 0.17 81.00 72.94 55.38 0.45 WizardLM-70b 49.51 52.70 0.27 83.30 80.68 71.93 0.58 Vicuna-33b 52.76 64.00 0.23 75.00 81.57 64.51 0.57 Alpaca-7b 30.58 37.92 0.15 23.60 43.31 33.79 0.16 Falcon-40b-instruct 24.54 67.50 0.14 80.00 76.60 56.70 0.47 MPT-30b-chat 30.70 50.40 0.20 24.53 87.12 70.73 0.38 Starchat 28.30 40.12 0.12 25.40 16.96 9.07 0.05 Pythia-12b 24.71 27.00 0.12 25.60 24.49 31.80 0.07
Appendix D Total Performance
Table 4 provides an aggregated view of the overall performance of each model across all benchmark datasets. The total performance scores were computed by averaging the normalized scores across the selected evaluation metrics, offering a holistic comparison of model capabilities.
| Benchmark | Model | Baseline | ||||
| MMLU-STEM | Llama-3.1-3B | SFT | 0.2331 | 0.2679 | 0.2434 | 0.2200 |
| DPO | 0.3130 | 0.4868 | 0.2528 | 0.3940 | ||
| Llama-3.1-8B | SFT | 0.2800 | 0.2460 | 0.2660 | 0.2240 | |
| DPO | 0.4902 | 0.4349 | 0.4867 | 0.2880 | ||
| Qwen2.5-3B | SFT | 0.2760 | 0.2200 | 0.2460 | 0.2500 | |
| DPO | 0.4706 | 0.4212 | 0.4400 | 0.4580 | ||
| Qwen2.5-7B | SFT | 0.2800 | 0.2280 | 0.2620 | 0.2780 | |
| DPO | 0.5120 | 0.3620 | 0.4860 | 0.2800 | ||
| MMLU-pro | Llama-3.1-3B | SFT | 0.1189 | 0.1230 | 0.1045 | 0.1332 |
| DPO | 0.1455 | 0.2193 | 0.1270 | 0.1516 | ||
| Llama-3.1-8B | SFT | 0.1004 | 0.1148 | 0.1045 | 0.1311 | |
| DPO | 0.1168 | 0.1025 | 0.1168 | 0.1107 | ||
| Qwen2.5-3B | SFT | 0.1025 | 0.0840 | 0.1230 | 0.0820 | |
| DPO | 0.2275 | 0.2254 | 0.2029 | 0.2111 | ||
| Qwen2.5-7B | SFT | 0.0861 | 0.1762 | 0.2275 | 0.1352 | |
| DPO | 0.2377 | 0.1721 | 0.2254 | 0.1700 | ||
| IFeval | Llama-3.1-3B | SFT | 0.2494 | 0.2410 | 0.2490 | 0.2206 |
| DPO | 0.3765 | 0.4940 | 0.4796 | 0.3033 | ||
| Llama-3.1-8B | SFT | 0.2494 | 0.4210 | 0.2470 | 0.2218 | |
| DPO | 0.2421 | 0.1882 | 0.2292 | 0.2494 | ||
| Qwen2.5-3B | SFT | 0.2190 | 0.2134 | 0.2122 | 0.2050 | |
| DPO | 0.3633 | 0.3058 | 0.3094 | 0.1715 | ||
| Qwen2.5-7B | SFT | 0.2290 | 0.2134 | 0.2122 | 0.2083 | |
| DPO | 0.3321 | 0.3177 | 0.3657 | 0.2407 | ||
| ARC-easy | Llama-3.1-3B | SFT | 0.2660 | 0.2460 | 0.2000 | 0.2340 |
| DPO | 0.6111 | 0.5547 | 0.2618 | 0.6679 | ||
| Llama-3.1-8B | SFT | 0.2180 | 0.2330 | 0.2400 | 0.2360 | |
| DPO | 0.2176 | 0.1540 | 0.4720 | 0.0680 | ||
| Qwen2.5-3B | SFT | 0.2410 | 0.2560 | 0.2480 | 0.2560 | |
| DPO | 0.8497 | 0.8754 | 0.5295 | 0.8157 | ||
| Qwen2.5-7B | SFT | 0.2550 | 0.2260 | 0.2280 | 0.2900 | |
| DPO | 0.5700 | 0.6359 | 0.6738 | 0.7134 | ||
| ARC-challenge | Llama-3.1-3B | SFT | 0.2556 | 0.2492 | 0.2266 | 0.2019 |
| DPO | 0.5122 | 0.4551 | 0.2466 | 0.5712 | ||
| Llama-3.1-8B | SFT | 0.2320 | 0.2297 | 0.2761 | 0.2268 | |
| DPO | 0.5463 | 0.1763 | 0.3596 | 0.2946 | ||
| Qwen2.5-3B | SFT | 0.2645 | 0.2483 | 0.2227 | 0.2343 | |
| DPO | 0.7398 | 0.8241 | 0.4470 | 0.7078 | ||
| Qwen2.5-7B | SFT | 0.2343 | 0.2552 | 0.2483 | 0.2390 | |
| DPO | 0.4432 | 0.5358 | 0.5847 | 0.3870 |